Please help me solve the following:
The perimeter of a rectangle is 122m. The length is 10m more than twice the width. Find the dimensions.

Answer by Theo(13342)

P = perimeter = 122 meters.
Perimeter = 2 * ( length + width )
Let:
length = L
width = W
This makes Perimeter = 2 * (L + W)
Since length = 2 * width, you get:
length = 2W
width = W
This makes Perimeter = 2 * 3W = 6W
Since P = 122 meters, you get:
122 meters = 6W
Divide both sides of the equation by 6 to get:
W = 20.3333333 which is the same as:
W = 20 and 1/3 meters.
Since L = 2W, then:
L = 40 and 2/3 meters.

Since P = 2 * (L + W) then:

P = 2 * (20 + 1/3 + 40 + 2/3) = 2 * 61 = 122 meters.

Your answer is:

W = 20 and 1/3 meters.
P = 40 and 2/3 meters.